# Time Series Forecasting with LLMs: Understanding and Enhancing Model Capabilities

### v1

@@ A.2.1 Real Datasets
We conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTraffic and TurkeyPower, as detailed in A.1. We apply the Seasonal-Trend decomposition using the LOESS (STL) technique, to decompose the original time series into trend, seasonal, and residual components. In those datasets, we obtain the periods through the nature of the data. For instance, the number of passengers is collected every month in AirPassengersDataset, and it’s natural to obtain that the period is 12. For the datasets without explicit periods, such as the IstanbulTraffic, the period was determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period. The strengths and the model performance can be seen in Table 1.
Subsequently, we compute trend strength and seasonal strength to measure all of those components. We use and MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 2). Notably, GPT-4 achieved a higher absolute PCC compared to GPT-3.5-turbo-instruct. It may be attributed to human feedback during GPT-4 training, as individuals may be more aware of seasonal and trend data. Interestingly, upon calculating the strengths of the output sequences, there is an increase in the seasonal component of the output generated by GPT-4, in contrast to the original test sequences. This may provide some insights for further research into the characteristics of the LLMs.
Following this, we conduct the counterfactual analysis with a systematic permutation to the input time series. We initiate by defining a sliding window that constitutes 10% of the total length of the time seriesm and add Gaussian noise into the data within this window data. Given the variability in the length of time sequences, we first scale the sequence through max-min normalization before incorporating it into the original sequence. The size of the sliding window, , is defined as: , where is the total length of the time series. We denote the start point of the window as and the Gaussian noise as , to the data within the sliding window, the modified data point, , is given by:
where , and is the data point after normalization. The change in Mean Squared Error () is defined as the metric to measure the sensitivity of the LLMs:
where as , reaches its maximum.
This method allows us to assess the impact of individual segments and thereby infer the interpretability of the time series segments that LLM predominantly focuses on. Our observations suggest that introducing noise towards the end of the time series significantly affects LLM’s performance, leading to the inference that LLM tends to give more weight to the latter part of the time series in most instances.
@@ A.2.2 Synthesized Datasets
To further validate our findings, we conducted experiments on synthesized datasets. We formulate the dataset through the expression , where ranges from 0 to 20, and the noise term follows a normal distribution . The coefficients and were uniformly sampled for 10 samples respectively, with and . Hence, both trend and seasonal strengths were carefully confined to ,, resulting in more discernible conclusions in the output. The chosen frequency, , was consistent across the experiments. We uniformly sampled 200 data points. Our observations validate that LLMs prefer input sequences with higher strengths and generate time series with higher seasonal strength. Furthermore, we observed that the model performance cannot be entirely explained by the linear combination of the strengths, implying that the influence of these strengths is not independent.
@@
To investigate the influence of the number of periods on model performance, we generated a dataset using the function , where denote the coefficients of the trend and seasonal components. We set , , and , uniformly sampled for 10 instances each, and and to represent the chosen frequencies. Similar to the previous experiments, ranges from 0 to 20 and follows the normal distribution . Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure6(a). This observation may be attributed to the LLMs’ challenge in recognizing and adapting to multiple periods, similar to human behavior.
A.3 Detail Prompts
Request output periodic prompts
Please output the period size of the predicted sequence as an integer. Output an integer directly.
In-context learning examples
For example, the sequence: 0.387952, 8.975192, 5.398713, -6.011139, -9.807413, -0.261663, 9.245404, 5.901009, -6.038525, -9.966693, -0.022652, 9.352823,5.317264, -5.809847, -9.565386, -0.325205, 8.711379, 6.176601, -5.911722, -9.078278.We can get the period number is 5. so, try to find out this period:
A.4 Figures and Tables
@@
| Sunspots | 7072.42 | 62.61 | 194.29 | 33094824.12 | 4209.81 | 17.69 | |
| GPT-3.5 | Wine | 24925885.81 | 3548.19 | 14.98 | 820422.05 | 760.03 | 16.51 |
| (gpt-3.5-turbo-1106) | Wooly | 955708.49 | 893.02 | 19.26 | 65.95 | 7.05 | 7.47 |
@@
| Sunspots | 4571.92 | 50.24 | 334.30 | 10105667.50 | 2438.74 | 9.85 | |
| GPT-4 | Wine | 14426570.88 | 2734.41 | 10.90 | 1211754.67 | 1017.44 | 21.85 |
| (gpt-4-turbo-preview) | Wooly | 1078968.96 | 959.42 | 20.41 | 85.22 | 6.89 | 7.77 |
@@
| Sunspots | 4000.19 | 46.45 | 138.69 | 11788774.81 | 2770.67 | 10.90 | |
| Llama-2 | Wine | 8286095.02 | 2261.30 | 8.97 | 636637.46 | 742.67 | 15.54 |
| (llama-2-13B) | Wooly | 389685.08 | 551.18 | 11.69 | 64.92 | 6.39 | 7.04 |
@@
| Wind | 1.78 | 0.79 | 116.31 | 36.78 | 4.02 | 12.38 | |
@@
| LLMTime with GPT-3.5-Turbo | 267.66 | 3.66 | 0.99 | |
| LLMTime with GPT-4-Turbo | 133.10 | 2.87 | 0.80 | |
@@
| LLMTime with GPT-3.5-Turbo | 598.45 | 5.81 | 1.36 | |
| LLMTime with GPT-4-Turbo | 661.80 | 7.24 | 1.63 | |
@@
| LLMTime with GPT-3.5-Turbo | 968.69 | 8.61 | 1.02 | |
| LLMTime with GPT-4-Turbo | 413.63 | 4.94 | 0.57 | |
@@
| LLMTime with GPT-3.5-Turbo | 251.61 | 4.27 | 20.42 | |
| LLMTime with GPT-4-Turbo | 194.52 | 5.30 | 16.10 |
| Dataset Name | Average Trend Strength | Average Seasonal Strength | Median Trend Strength | Median Seasonal Strength | Test Set Trend Strength | Test Set Seasonal Strength |
| AirPassengersDataset | ||||||
| AusBeerDataset | ||||||
| GasRateCO2Dataset | ||||||
| MonthlyMilkDataset | ||||||
| SunspotsDataset | ||||||
| WineDataset | ||||||
| WoolyDataset | ||||||
| HeartRateDataset | ||||||
| IstanbulTraffic | ||||||
| TurkeyPower |
| Dataset Name | Average Trend Strength | Average Seasonal Strength | Median Trend Strength | Median Seasonal Strength | Test Set Trend Strength | Test Set Seasonal Strength |
| AirPassengersDataset | ||||||
| AusBeerDataset | ||||||
| GasRateCO2Dataset | ||||||
| MonthlyMilkDataset | ||||||
| SunspotsDataset | ||||||
| WineDataset | ||||||
| WoolyDataset | ||||||
| HeartRateDataset | ||||||
| IstanbulTraffic | ||||||
| TurkeyPower |

### v3

@@ A.2.1 Real Datasets
We begin by comparing the performance of Large Language Models (LLMs) with traditional time series forecasting methods. The results are depicted in Figures Figure 4 and Figure 5, while Tables Table 9 and Table 10 present the computed metrics in tabular format.
To understand the preferences of LLMs, we conducted experiments on ten commonly used datasets: HeartRateDataset, GasRateCO2Dataset, AirPassengersDataset, AusBeerDataset, MonthlyMilkDataset, SunspotsDataset, WineDataset, WoolyDataset, IstanbulTrafficDataset and TurkeyPowerDataset. We apply the Seasonal-Trend decomposition using the LOESS (STL) technique to decompose the original time series into trend, seasonal, and residual components. In those datasets, we obtain the periods through the nature of the data. For instance, the number of passengers is collected monthly in AirPassengersDataset, and it’s natural to obtain that the period is 12. For the datasets without explicit periods, such as the IstanbulTrafficDataset, the period is determined through the periodogram, a widely used tool in signal processing assisting the identification of the time series period. The strengths and the model performance can be seen in Table 6.
Subsequently, we compute trend strength and seasonal strength to measure all of those components. We use MAPE to compute the Pearson correlation coefficients (PCC) across every two indexes and observe a relatively strong correlation between the strengths and model performance, signifying that LLMs perform better when the input time series possesses higher trend and seasonal strength (Shown in Table 1). Notably, GPT-4-turbo achieved a higher absolute PCC compared to GPT-3.5-turbo. It may be attributed to human feedback during GPT-4-turbo training, as individuals may be more aware of seasonal and trend data. This may provide some insights for further research into the characteristics of the LLMs with time series forecasting.
We also conduct a counterfactual analysis using a systematic permutation of the input time series. We begin by defining a sliding window that determines the length of the periods and add Gaussian noise to the data within this window. To reduce costs, we move the window by the period length. This method allows us to assess the importance of segments that LLM predominantly focuses on. Our observations suggest that introducing noise towards the end of the time series significantly affects LLM’s performance, leading to the inference that LLM tends to give more weight to the latter part of the time series in most instances.
@@ A.2.2 Synthesized Datasets
To investigate the influence of the number of periods on model performance, we generated a dataset using the function , where denote the coefficients of the trend and seasonal components. We set , , and , uniformly sampled for 10 instances each, and and to represent the chosen frequencies. Similar to the previous experiments, ranges from 0 to 20 and follows the normal distribution . Our results reveal that LLMs exhibit worse performance when input sequences contain multiple periods, even when the seasonal strength is carefully controlled to be nearly unchanged, as is shown in Figure 8. This observation may be attributed to the LLMs’ challenge in recognizing and adapting to multiple periods, similar to human behavior.
A.3 Detailed Prompts for the Period Prediction
GPT-3.5-turbo and GPT-4-turbo.
You are a helpful assistant that specializes in time series analysis. The user will provide a sequence. The sequence is separated by commas. You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence. Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence. The underlying cycle of the sequence is strictly below 8. Please infer the cycle without producing any additional text. Again, the sequence is separated by commas. Sequence:
Gemini.
You are a helpful assistant that specializes in time series analysis. The user will provide a sequence. The sequence is represented by decimal strings separated by commas. You need to infer the most probable underlying cycle of the sequence, even though there may also be a trend in this sequence. Do not assume that the underlying cycle has to stick to the popular cycles like 7 (days in a week), 12 (months in a year) and 30 (days in a month), just infer the cycle fully based on the inherent cycle of the given sequence. The underlying cycle of the sequence is strictly below 8. Please infer the cycle without producing any additional text. Sequence:
A.4 Tests on Data Leakage
Since the detailed datasets used for training GPT-4 and Gemini are not directly accessible, we conducted several experiments to indirectly investigate the data leakage issue. First, we performed an "Acknowledge Test" where we asked the LLMs if they recognize the dataset based on its name. Next, we conducted a "Series Test" by asking the LLMs to predict the first 20 steps of the datasets. Finally, we carried out a "Dataset Detection" test, where we fed the first 20 steps into the LLMs and asked them to identify the dataset. All these experiments help us determine whether the LLMs only recognize the name or have detailed knowledge of the datasets.
As shown in 4, we observe that both GPT and Gemini are generally aware of the content of most datasets. However, detailed sequence data is primarily known only for the AirPassengers dataset when using GPT, and this level of detail is not known for other datasets. During the experiments on the Dataset Detection, it is found that both GPT and Gemini could identify which dataset is being used based on the first 20 data points of the time series. All these results indicate that while they can recognize and identify datasets from limited information, they typically do not possess detailed knowledge of the sequence data for a broader range of datasets.
| Datasets | Acknowledge Test (GPT) | Acknowledge Test (Gemini) | Series Test (GPT) | Series Test (Gemini) | Dataset Detection (GPT) | Dataset Detection (Gemini) |
| AirPassengers | Yes | Yes | Yes | No | No | No |
| AusBeer | No | Yes | No | No | No | No |
| GasRateCO2 | No | Yes | No | No | No | No |
| MonthlyMilk | Yes | Yes | No | No | No | No |
| Sunspots | Yes | Yes | No | No | No | No |
| Wine | Yes | Yes | No | No | No | No |
| Wooly | No | No | No | No | No | No |
| HeartRate | Yes | Yes | No | No | No | No |
A.5 Computational Cost
We list the average token length cost associated with external knowledge enhancing and natural language paraphrasing for reference. Avg Token Length(ori) is the prompt Length of the unexecuted method, and Avg Token Length(EKE, NLP) is the prompt length after executing the corresponding policy. It is noted that Natural Language Paraphrasing is judged one by one through hard coding. Besides, there is a length check after transformation, so it is guaranteed that a certain length can be obtained each time. The results are shown in 5.
| Datasets | Avg Token Length (ori) | Avg Token Length (EKE) | Avg Token Length (NLP) |
| AirPassengers | 200 | 224 | 797 |
| AusBeer | 200 | 220 | 797 |
| GasRateCO2 | 200 | 211 | 797 |
| MonthlyMilk | 200 | 218 | 797 |
| Sunspots | 200 | 217 | 797 |
| Wine | 200 | 217 | 797 |
| Wooly | 200 | 216 | 797 |
| HeartRate | 200 | 214 | 797 |
| Dataset Name | GPT4-MAPE | GPT3.5-MAPE | Trend Strength | Seasonal Strength |
| AirPassengersDataset | 6.80 | 9.98 | 1.00 | 0.98 |
| AusBeerDataset | 3.69 | 5.12 | 0.99 | 0.96 |
| MonthlyMilkDataset | 5.12 | 6.25 | 1.00 | 0.99 |
| SunspotsDataset | 334.30 | 194.29 | 0.81 | 0.28 |
| WineDataset | 10.90 | 14.98 | 0.67 | 0.92 |
| WoolyDataset | 20.41 | 19.26 | 0.96 | 0.82 |
| IstanbulTrafficGPT | 47.29 | 60.11 | 0.31 | 0.72 |
| GasRateCO2Dataset | 4.21 | 5.97 | 0.65 | 0.50 |
| HeartRateDataset | 7.90 | 6.75 | 0.42 | 0.49 |
| TurkeyPower | 3.36 | 3.52 | 0.90 | 0.88 |
A.6 Figures and Tables
@@
| MSE | MAE | MAPE | MSE | MAE | MAPE | ||
| AusBeer | 598.45 | 5.81 | 1.36 | 841.68 | 23.59 | 5.62 | |
| GasRateCO2 | 3.16 | 0.46 | 0.85 | 10.88 | 2.66 | 4.73 | |
| MonthlyMilk | 968.69 | 8.61 | 1.02 | 7507.13 | 66.28 | 112.77 | |
| Sunspots | 251.61 | 4.27 | 20.42 | 6556.55 | 58.95 | 217.94 | |
| (GPT-3.5-turbo-1106) | Wine | 11403.89 | 96.95 | 37.04 | 30488.60 | 388.28 | 15.83 |
| Wooly | 12110.16 | 33.23 | 4.07 | 526903.08 | 574.58 | 12.00 | |
| HeartRate | 4.38 | 0.55 | 0.57 | 76.83 | 7.15 | 7.42 | |
| Istanbul-Traffic | 224.17 | 3.74 | 8.81 | 335.05 | 6.75 | 11.68 | |
| Turkey Power | 24382136.98 | 1843.64 | 4.68 | 3882704.14 | 1315.6 | 3.58 | |
| ETTh1 | 1.21 | 0.48 | 54.17 | 5.64 | 2.71 | 1.625 | |
| ETTm2 | 0.81 | 0.36 | 27.33 | 3.46 | 2.17 | 1.178 | |
| US Births | 926136.72 | 633.37 | 7.67 | 1323052.46 | 904.81 | 9.61 | |
| Saugeen River Flow | 2870.19 | 18.31 | 19.43 | 4100.27 | 26.16 | 27.76 | |
| AusBeer | 661.80 | 7.24 | 1.63 | 513.49 | 18.57 | 4.28 | |
| GasRateCO2 | 2.28 | 0.41 | 0.75 | 7.27 | 2.32 | 4.18 | |
| MonthlyMilk | 413.63 | 4.94 | 0.57 | 4442.18 | 50.75 | 172.82 | |
| Sunspots | 194.52 | 5.30 | 16.10 | 3374.70 | 41.87 | 321.11 | |
| (GPT-4-turbo-preview) | Wine | 56138.87 | 54.67 | 23.63 | 22488.17 | 253.08 | 9.98 |
| Wooly | 18063.64 | 11.06 | 25.06 | 942987.19 | 871.64 | 18.55 | |
| HeartRate | 11.64 | 1.21 | 1.30 | 988.14 | 26.57 | 29.22 | |
| Istanbul-Traffic | 176.91 | 3.88 | 9.67 | 195.33 | 5.53 | 10.03 | |
| Turkey Power | 60601807.53 | 3118.43 | 8.03 | 113873.28 | 814.46 | 2.17 | |
| ETTh1 | 1.20 | 0.49 | 47.62 | 4.73 | 1.53 | 3.282 | |
| ETTm2 | 0.45 | 0.27 | 23.62 | 2.30 | 1.034 | 1.607 | |
| US Births | 676264.81 | 501.13 | 4.81 | 966092.59 | 678.55 | 7.12 | |
| Saugeen River Flow | 3100.57 | 19.18 | 21.86 | 4190.05 | 27.41 | 32.15 | |
| Llama-2 | AirPassengers | 751.34 | 6.77 | 1.53 | 1317.9 | 55.49 | 11.18 |
| AusBeer | 591.75 | 23.25 | 5.41 | 644.82 | 17.88 | 4.08 | |
| GasRateCO2 | 10.16 | 2.89 | 5.16 | 12.78 | 2.97 | 5.47 | |
| MonthlyMilk | 851.17 | 84.83 | 9.46 | 3410.20 | 41.40 | 240.25 | |
| Sunspots | 1483.29 | 33.27 | 17.79 | 4467.67 | 48.95 | 91.79 | |
| (llama-2-13B) | Wine | 102434.52 | 852.97 | 34.72 | 951194.94 | 240.08 | 9.45 |
| Wooly | 12180.05 | 83.99 | 16.92 | 675062.52 | 736.04 | 15.83 | |
| HeartRate | 49.8 | 5.84 | 6.53 | 75.58 | 7.11 | 7.94 | |
| Istanbul-Traffic | 306.80 | 5.39 | 7.24 | 438.28 | 7.28 | 9.81 | |
| Turkey Power | 3278744.18 | 2191.34 | 28.76 | 2919773.15 | 1388.10 | 3.70 | |
| ETTh1 | 1.47 | 0.87 | 58.34 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 0.84 | 0.41 | 29.86 | 3.31 | 2.07 | 2.153 | |
| US Births | - | - | - | - | - | - | |
| Saugeen River Flow | - | - | - | - | - | - | |
| Gemini-Pro-1.0 | AirPassengers | 4474.54 | 31.54 | 7.02 | 6392.21 | 63.57 | 14.03 |
| AusBeer | 278.45 | 10.05 | 2.29 | 397.78 | 14.36 | 3.27 | |
| GasRateCO2 | 13.29 | 2.50 | 4.38 | 18.99 | 3.57 | 6.46 | |
| MonthlyMilk | 440.29 | 11.91 | 1.39 | 628.98 | 17.01 | 1.99 | |
| Sunspots | 438.29 | 10.47 | 1.21 | 626.03 | 14.94 | 1.73 | |
| (gemini-1.0-pro) | Wine | 181008.34 | 2557.62 | 10.89 | 258584.78 | 3645.23 | 14.60 |
| Wooly | 45.44 | 4.47 | 4.93 | 64.92 | 6.39 | 7.04 | |
| HeartRate | 40.57 | 4.20 | 4.67 | 57.96 | 6.01 | 6.66 | |
| Istanbul-Traffic | 267.43 | 5.69 | 8.37 | 321.56 | 7.32 | 9.71 | |
| Turkey Power | 45674.41 | 2973.54 | 11.21 | 103203.37 | 2195.68 | 6.07 | |
| ETTh1 | 1.17 | 0.74 | 54.86 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 0.88 | 0.39 | 21.82 | 3.31 | 2.07 | 2.153 | |
| US Births | 467743.19 | 440.68 | 4.27 | 687862.05 | 595.51 | 6.28 | |
| Saugeen River Flow | 2857.23 | 19.34 | 18.80 | 4081.75 | 25.91 | 26.87 | |
@@
| Sunspots | 7072.42 | 62.61 | 194.29 | 6556.55 | 58.95 | 217.94 | |
| GPT-3.5-turbo-1106 | Wine | 24925885.81 | 3548.19 | 14.98 | 30488.60 | 388.28 | 15.83 |
| (GPT-3.5-turbo-1106) | Wooly | 955708.49 | 893.02 | 19.26 | 526903.08 | 574.58 | 12.00 |
| HeartRate | 59.83 | 6.44 | 6.75 | 76.83 | 7.15 | 7.42 | |
@@
| Sunspots | 4571.92 | 50.24 | 334.30 | 3374.70 | 41.87 | 321.11 | |
| GPT-4-turbo-preview | Wine | 14426570.88 | 2734.41 | 10.90 | 22488.17 | 253.08 | 9.98 |
| (GPT-4-turbo-preview) | Wooly | 1078968.96 | 959.42 | 20.41 | 942987.19 | 871.64 | 18.55 |
| HeartRate | 78.99 | 6.96 | 7.90 | 988.14 | 26.57 | 29.22 | |
@@
| Sunspots | 4000.19 | 46.45 | 138.69 | 4467.67 | 48.95 | 91.79 | |
| Llama-2 | Wine | 8286095.02 | 2261.30 | 8.97 | 951194.94 | 240.08 | 9.45 |
| (llama-2-13B) | Wooly | 389685.08 | 551.18 | 11.69 | 675062.52 | 736.04 | 15.83 |
| HeartRate | 112.17 | 7.86 | 8.93 | 75.58 | 7.11 | 7.94 | |
@@
| AirPassengers | 5237.85 | 51.92 | 11.08 | 6392.21 | 63.57 | 14.03 | |
| AusBeer | 325.45 | 10.84 | 1.86 | 397.78 | 14.36 | 3.27 | |
| GasRateCO2 | 15.54 | 3.23 | 4.43 | 18.99 | 3.57 | 6.46 | |
| MonthlyMilk | 491.26 | 15.18 | 1.13 | 628.98 | 17.01 | 1.99 | |
| Sunspots | 491.64 | 11.15 | 1.27 | 626.03 | 14.94 | 1.73 | |
| Gemini-1.0-pro | Wine | 210818.24 | 3230.41 | 8.35 | 258584.78 | 3645.23 | 14.60 |
| (gemini-1.0-pro) | Wooly | 51.04 | 5.70 | 7.93 | 64.92 | 6.39 | 7.04 |
| HeartRate | 47.45 | 4.83 | 4.67 | 57.96 | 6.01 | 6.66 | |
| Istanbul-Traffic | 1253.74 | 28.25 | 5.42 | 1531.37 | 34.74 | 7.42 | |
| TSMC-Stock | 153.73 | 5.02 | 1.05 | 188.18 | 6.67 | 1.65 | |
| Turkey Power | 83522.95 | 1812.31 | 5.51 | 103203.37 | 2195.68 | 6.07 | |
| ETTh1 | 2.92 | 1.45 | 2.88 | 4.84 | 1.79 | 3.178 | |
| ETTm2 | 2.00 | 1.74 | 1.22 | 3.31 | 2.07 | 2.153 | |
@@
| N-Beats | 3994.55 | 54.95 | 12.81 | |
| DeepAR | 184222.64 | 421.99 | 98.42 | |
| Prophet | 7345.31 | 43.87 | 8.62 | |
| LLMTime with GPT-3.5-Turbo | 6244.07 | 61.39 | 14.43 | |
| LLMTime with GPT-4-Turbo | 1317.9 | 55.49 | 11.18 | |
| LLMTime with Gemini-1.0-pro | 6392.21 | 63.57 | 14.03 | |
| LLMtime with Llama-2 | 1286.25 | 28.04 | 6.07 | |
@@
| N-Beats | 250.61 | 14.42 | 3.53 | |
| DeepAR | 16197.17 | 40.23 | 9.89 | |
| Prophet | 6323.89 | 28.76 | 6.92 | |
| LLMTime with GPT-3.5-Turbo | 841.68 | 23.59 | 5.62 | |
| LLMTime with GPT-4-Turbo | 513.49 | 18.57 | 4.28 | |
| LLMTime with Gemini-1.0-pro | 397.78 | 14.36 | 3.27 | |
| LLMtime with Llama-2 | 644.82 | 17.88 | 4.08 | |
@@
| N-Beats | 3140.89 | 51.57 | 6.07 | |
| DeepAR | 728289.50 | 851.30 | 99.22 | |
| Prophet | 663.41 | 25.76 | 2.92 | |
| LLMTime with GPT-3.5-Turbo | 7507.13 | 66.28 | 112.77 | |
| LLMTime with GPT-4-Turbo | 4442.18 | 50.75 | 172.82 | |
| LLMTime with Gemini-1.0-pro | 628.98 | 17.01 | 1.99 | |
| LLMtime with Llama-2 | 3410.20 | 41.40 | 240.25 | |
@@
| N-Beats | 4877.59 | 56.58 | 105.55 | |
| DeepAR | 3421.02 | 48.93 | 132.76 | |
| Prophet | 6303.57 | 76.83 | 67.97 | |
| LLMTime with GPT-3.5-Turbo | 6556.55 | 58.95 | 217.94 | |
| LLMTime with GPT-4-Turbo | 3374.70 | 41.87 | 321.11 | |
| LLMTime with Gemini-1.0-pro | 626.03 | 14.94 | 1.73 | |
| LLMtime with Llama-2 | 4467.67 | 48.95 | 91.79 |
| Dataset | Method | MSE | MAE | MAPE |
| WineDataset | Exponential Smoothing | 23709576.52 | 3370.78 | 14.23 |
| SARIMA | 1150166.94 | 966.57 | 20.76 | |
| Cyclical Regression | 7873785.27 | 2148.24 | 8.52 | |
| AutoARIMA | 698661.90 | 646.03 | 14.07 | |
| FFT | 1031170.45 | 867.83 | 18.60 | |
| StatsForecastAutoARIMA | 20040877.37 | 2853.17 | 12.05 | |
| Naive Mean | 11557786.19 | 2200.04 | 8.80 | |
| Naive Seasonal | 879447.22 | 724.23 | 15.52 | |
| Naive Drift | 9609576.04 | 1833.38 | 7.36 | |
| Naive Moving Average | 9070696.99 | 1719.17 | 6.90 | |
| N-Beats | 5418377.00 | 1887.30 | 7.68 | |
| DeepAR | 715027008.00 | 26236.14 | 89.91 | |
| Prophet | 4846922.27 | 2201.57 | 8.27 | |
| LLMTime with GPT-3.5-Turbo | 30488.60 | 388.28 | 15.83 | |
| LLMTime with GPT-4-Turbo | 22488.17 | 253.08 | 9.98 | |
| LLMTime with Gemini-1.0-pro | 258584.78 | 3645.23 | 14.60 | |
| LLMtime with Llama-2 | 951194.94 | 240.08 | 9.45 | |
| WoolyDataset | Exponential Smoothing | 24925885.81 | 3548.19 | 14.98 |
| SARIMA | 812352.21 | 759.07 | 16.37 | |
| Cyclical Regression | 1032574.82 | 962.72 | 22.14 | |
| AutoARIMA | 838852.91 | 786.25 | 16.84 | |
| FFT | 1012255.35 | 945.20 | 20.80 | |
| StatsForecastAutoARIMA | 917617.19 | 858.57 | 18.91 | |
| Naive Mean | 816762.31 | 764.73 | 16.12 | |
| Naive Seasonal | 1051110.81 | 982.25 | 22.19 | |
| Naive Drift | 812352.21 | 759.07 | 16.37 | |
| Naive Moving Average | 1032574.82 | 962.72 | 22.14 | |
| N-Beats | 653104.31 | 743.54 | 15.96 | |
| DeepAR | 243831.14 | 4897.85 | 94.89 | |
| Prophet | 365241.98 | 891.70 | 34.65 | |
| LLMTime with GPT-3.5-Turbo | 526903.08 | 574.58 | 12.00 | |
| LLMTime with GPT-4-Turbo | 942987.19 | 871.64 | 18.55 | |
| LLMTime with Gemini-1.0-pro | 64.92 | 6.39 | 7.04 | |
| LLMtime with Llama-2 | 675062.52 | 736.04 | 15.83 | |
| HeartRateDataset | Exponential Smoothing | 11.16 | 1.38 | 1.49 |
| SARIMA | 12.98 | 1.34 | 1.61 | |
| Cyclical Regression | 13.58 | 1.31 | 1.20 | |
| AutoARIMA | 13.26 | 1.25 | 1.39 | |
| FFT | 13.95 | 1.16 | 1.34 | |
| StatsForecastAutoARIMA | 10.53 | 1.27 | 1.39 | |
| Naive Mean | 12.02 | 1.27 | 1.26 | |
| Naive Seasonal | 10.55 | 1.32 | 1.31 | |
| Naive Drift | 10.60 | 1.15 | 1.30 | |
| Naive Moving Average | 12.13 | 1.27 | 1.34 | |
| N-Beats | 72.11 | 7.10 | 7.40 | |
| DeepAR | 286.82 | 15.67 | 16.36 | |
| Prophet | 88.93 | 10.97 | 6.54 | |
| LLMTime with GPT-3.5-Turbo | 76.83 | 7.15 | 7.42 | |
| LLMTime with GPT-4-Turbo | 988.14 | 26.57 | 29.22 | |
| LLMTime with Gemini-1.0-pro | 57.96 | 6.01 | 6.66 | |
| LLMtime with Llama-2 | 75.58 | 7.11 | 7.94 | |
| Weather | Exponential Smoothing | 1684.38 | 31.60 | 6.79 |
| SARIMA | 1943.81 | 33.33 | 7.09 | |
| Cyclical Regression | 1700.73 | 30.77 | 7.15 | |
| AutoARIMA | 7315.10 | 57.44 | 11.70 | |
| FFT | 2752.02 | 38.90 | 8.87 | |
| StatsForecastAutoARIMA | 2479.55 | 38.06 | 8.16 | |
| Naive Mean | 39879.84 | 168.27 | 36.44 | |
| Naive Seasonal | 5057.47 | 52.81 | 11.89 | |
| Naive Drift | 5466.23 | 60.58 | 14.70 | |
| Naive Moving Average | 5057.47 | 52.81 | 11.89 | |
| N-Beats | 4532.84 | 39.21 | 23.49 | |
| DeepAR | 6325.75 | 35.97 | 16.59 | |
| Prophet | 3768.15 | 29.36 | 24.01 | |
| LLMTime with GPT-3.5-Turbo | 224.54 | 3.07 | 0.83 | |
| LLMTime with GPT-4-Turbo | 111.65 | 2.40 | 0.64 | |
| LLMTime with Gemini-1.0-pro | 176.32 | 3.72 | 0.75 | |
| LLMtime with Llama-2 | 215.39 | 4.07 | 1.31 |